OpenStudy (anonymous):

find the graph of the inequality y<-1/5x+1

OpenStudy (anonymous):

@jim_thompson5910 can you help me with this one too? :)

OpenStudy (anonymous):

@Jhannybean can you maybe try and help me out too?

OpenStudy (jhannybean):

\[y<-\frac{1}{5}x+1\]First draw out the line \(y=-\dfrac{1}{5}x+1\) 

OpenStudy (anonymous):

would it be a solid line or dashed?

OpenStudy (jhannybean):

Now above the line means \(y > -\dfrac{1}{5}x+1\) Below the line means \(y < -\frac{1}{5}x+1\) Oh yes, and the line would be dashed, because it maynot inslude vaues associated with the line \(y=-\dfrac{1}{5}x+1\)

OpenStudy (jhannybean):